# Physics-Informed Multi-Agent Reinforcement Learning for Distributed Multi-Robot Problems

### v1

@@ VII Conclusions
We have conducted ablation studies and comparative simulations against other state-of-the-art multi-agent reinforcement learning approaches in six scenarios, covering a wide variety of cooperative and competitive behaviors such as collision avoidance, navigation, transport, evasion and monitoring. In all the cases, our proposed approach exhibits superior performance in terms of cumulative reward per robot and scalability. Our approach, without further training, scales and achieves the same performance of other methods that are trained, ad hoc, with a specific number of robots. Nevertheless, there is still room for improvement in terms of generalization with respect to environmental conditions such as stage size. These results lead to the following conclusion: the combination of physics-informed neural networks and reinforcement learning techniques is a promising research line to address multi-robot problems.
@@ Appendix B Soft Actor-Critic Hyperparameters
The following table details the parameterization of the soft actor-critic algorithm for the different scenarios. {longtblr}[ caption = Soft actor-critic hyperparameters., entry = Short Caption, label = tab:sac_parameters ]colsep = 2pt, colspec = ||c|c|c|| Parameter Scenario Value
optimizer all Adam [79]
$r_{comm}$pbox4.5cmreverse transport, navigation $0.45$m
sampling $0.75$m
@@
adversarial $6\times 10^{5}$
[…]
gradients
all False
[…]
scaling
all False

### v3

@@ VII Conclusions
We conducted ablation studies and simulations in comparison to state-of-the-art multi-agent reinforcement learning approaches in seven scenarios, covering a wide variety of cooperative and competitive behaviors such as collision avoidance, navigation, transport, evasion and monitoring, and including a validation experiment in a realistic robotic platform. In all the cases, our proposed approach exhibits superior performance in terms of cumulative reward per robot and scalability. Our approach, without further training, scales and achieves the same performance of other methods that are trained, ad hoc, with a specific number of robots. Extensive validation in a real multi-robot navigation setting demonstrated that policies trained in an ideal simulator exhibit great zero-shot sim-to-real transfer capabilities, even under imperfect communication. Nevertheless, there is still room for improvement in terms of generalization with respect to environmental conditions such as stage size. These results lead to the following conclusion: the combination of physics-informed neural networks and reinforcement learning techniques is a promising research line to address multi-robot problems.
A important future direction is to utilize the physics-informed port-Hamiltonian parameterization of the control policies to provide performance or safety guarantees. For instance, passivity theory or control barrier functions can be employed to ensure convergence properties of the control policy or design safety modules on top of the control policy that provide collision avoidance guarantees. It also remains to be explored how to bypass the three-message passing scheme to compute $\partial H_{{\bm{\theta}}}(\mathbf{x})/\partial\mathbf{x}_{j}$ and rely only on single-message broadcasting protocols or perception cues since it is currently a computation and implementation bottleneck.
@@ Appendix B Soft Actor-Critic Hyperparameters
[…]